9) A space probe, designed to land on a distant planet, is equipped to obtain samples of material, identify them according to general molecular type, and analyze their chemical makeup. Results obtained from samples from one planet indicate that the nucleic acid molecules contain just three types of nucleotides. Polypeptide molecules from this planet are found to be made of 26 different kinds of amino acids.

a) What is the minimal number of bases required in a codon in order to code for the 26 kinds of amino acids.

b) If, in addition, the code were to include two different terminator codons, could the codons be of the size designated in you answer to a)? Explain.
Answer #1

a)There are 26 amino acids in the polypeptide. One codon encodes one amino acid. One codon consists of three bases. That's why (26×3)+3= 78+3=81 bases are responsible for synthesis of this polypeptide. Here +3 for stop codon.

b)If there are two terminator codons in the codes of this polypeptide, then the amino acids number is different in the polypeptide. The codon number also changed because the presence of additional stop codon leads to stop the translation process before it acquired the proper chain. That's why the codes should be altered for proper polypeptide.